OpenStudy (anonymous):

One number is three less than four times another. If their difference is 48, what is the larger number?

OpenStudy (bahrom7893):

x = 4y-3 x-y = 48

OpenStudy (bahrom7893):

solve the system..

OpenStudy (bahrom7893):

come on you can do this.. plug in the x: 4y-3-y = 48 3y-3=48 y-1=18 y = 19

OpenStudy (bahrom7893):

x-19=48 x = 48+19 x = 67 <,larger number